The Mixing Department of Crown Cement has the following cost and production data for



the month of June. [10]



Costs:



Work in process, June 1



Direct materials: 100% complete



Conversion costs: 20% Complete



$90,000



60,000



Cost of work in process, June 1 150,000



Costs incurred during production in June



Direct materials



Conversion costs



$600,000



280,000



Costs incurred in April 880,000



Units transferred put totaled 15,000. Ending work in process was 1200that are 100%



complete as to material and 40% complete as to conversion costs.



Instructions



a. Compute the equivalent units of production for (1) materials and (2)



conversion costs for the month of June. [4]



b. Compute the unit costs for the month. [3]



c. Determine the costs to be assigned to the units transferred out and in ending



work in process.

1
Expert's answer
2022-02-20T16:18:36-0500

a.

"equivalent units materials=1\times600 000=600 000"

"conversion costs=280 000\times0.4=112000"

b.

"\frac{600 000}{15000}=40"


"\frac{112000}{1200}=93.33"

c.

"150 000+1\times600 000+280 000\times0.4=862 000"

put totaled=15000+1200=16200